OpenStudy (tmagloire1):

Determine the equation of a line, in slope-intercept form, that passes through the points (9, –2) and (12, 0).

OpenStudy (anonymous):

m(x-x0)=y-y0 m(3)=2 3m=2 m=2/3 y=2x/3-8

OpenStudy (tmagloire1):

Thank you so much!

OpenStudy (anonymous):

np